If an analog temperature sensor is used with an 8 bit ADC (analog to digital converter) and a 5 volt power source, what is the smallest temperature increment you could measure? That is, what is the resolution here?
Answer #1

The smallest tempreture increment is given by resolution of a ADC

total number of bits used here = 8 bits

power supply = 5Volts

total number of steps= 28 - 1= 256-1=255

step size = resoution of ADC= (power voltage)/(total nomber of steps)

resolution= 5 volts/255= 0.01960Volts= 19.60 mV

smallest tempreture increment corresponding to 19.60mV is the smallest tempreture that can be measured.